Which of the following compounds will violate the octet rule? Answer A:03 Answer B: NCI3 Answer C: CO2 Answer D: NO2
Answer #1

NO₂ 0 – 0=0 - N c: CO2 o = c=0 NO₂ 30-N=ö Elections around N=7 NOviolates octet sule.According to octet rule , there must be 8 electrons in the valence shell or an atom.

In the Lewis structure of a compound there must be 8 electrons around each atom.

Steps to draw Lewis structure :- 1. Write the chemical symbols of each element involved in the structure. 2. Count the number